# Heavy-element damage seeding in proteins under XFEL illumination

### v4

@@ IV Damage by different trace elements
In this section, we present the results of simulations designed to examine how the energy-level structure of a trace, heavy element in a biological target affects electronic damage. Ionization dynamics were modeled in targets with the atomic ratio of non-derivative lysozyme, but with sulfur (1% of the atoms) replaced by ‘dopant’ atoms of some element. The ‘N-doped’ and ‘S-doped’ targets considered in this section thus correspond, respectively, to the light-atom control and non-derivative lysozyme targets considered in Sec. III.
The total number of modeled atomic configurations scales like the factorial of the total number of subshells of an element’s ground state, making heavier elements very costly to simulate. For computational expedience, shells of elements were modeled with a single angular momentum (i.e. a single energy level). Benchmarks showed this approximation had negligible consequences for the overall dynamics (see Fig. S2).
@@ IV.1 Photon and primary electron energy
Fig. 7(a) shows that the choice of dopant substantially impacts . At all energies, sees a wide range of values across the targets. Most traces show decreasing damage with increasing photon energy, consistent with prior studies [2]; however, the traces for the targets doped with Fe2+ (=26), Zn2+ (=30), and Se (=34) buck this trend, with the transition over the K-edge of the dopant roughly doubling in each case. This reflects the substantial contribution of the K-shell to the primary ionization of these dopants; the 1s orbitals dominate the photoabsorption cross-sections of their atoms and are replenished on a subfemtosecond timescale by Auger cycling (see Sec. III.1).
@@
Perhaps more surprisingly, continues to increase above the K-edges for 2 despite decreases in the photoabsorption cross-section. To understand this, one can consider a free electron of energy interacting with a gas of hydrogenic atoms, with electrons bound by energy . As falls, the EII cross-section grows, causing an increased ionization rate down to [67, 75]. Fig. 8 sketches typical showers of secondary electrons from a single primary electron in the regime. It can be seen that the increase in ionization cross-section as the impacting electron’s velocity falls creates a trade-off: higher-energy cascades progress more slowly, but they have the potential to free a greater number of electrons [3]. When counting all ionization events over a fixed duration, there will be a maximally ionizing cascade energy. This implies the existence of a ‘maximally damaging’ cascade energy , though of lower energy since the earlier ionizations of a more rapid cascade have a more significant influence on the scattering pattern (further magnified by Bragg gating [76, 77, 3]). Due to the higher photoabsorption cross-section near the edge, the most damaging photons eject primary electrons with energies below ; however, simulations that control for the rate of primary ionization show the difference to be slight (about 0.5 keV, see Appendix C).
@@
When plotted against the lowest photoelectron energy (LPE) at emission instead of the photon energy, the traces form two distinct groups: the ‘K-group’, targets doped by an element whose ‘deepest ionizable shell’ (DIS) in the considered energy range is the K-shell (N, S, Fe, Zn, Se), and the ‘L-group’, where it is the L-shell (Ag, Xe). The average ionization of carbon atoms in the sample follows a similar trend (Fig. 7(c)). The split in groups again appears to be a result of Auger cycling. As all heavy dopants’ DIS are maintained near maximum occupancy by decay processes, the production rate of ionizing electrons by Ag+ and Xe8+ is scaled by a factor of 4 relative to the other dopants at an equivalent LPE, due to the higher electron capacity of their L-shells.
@@
In general, the variation in electronic damage across the targets with respect to photon frequency is predominantly controlled by the LPE of the dopants and the shell number of the DIS. However, the traces of the N, S, and Fe-doped targets deviate from the main K-group in Fig. 7(b) and (c) for photon energies at or below 8 keV. In these cases, the photoelectrons ejected from the light atoms are of a low enough energy to contribute significantly to their secondary ionization (relative to the dopants). Since the Zn and Se K-edges are above 8 keV, their traces show no deviation. It is unsurprising that the L-group does not contain such outliers, as the primary electron contribution of the L-group dopants is much more dominant.
@@ IV.2 Damage landscape
Contour plots of the charge accumulated by the carbon atoms, as functions of photon energy and fluence, (Fig. 9(a)) or pulse width (Fig. 9(b)), show broad and significant differences due to the dopants. The effect of sulfur on the dynamics is relatively consistent; across the parameter space of Fig. 9, the S-doped target experiences levels of ionization that the pure CNO target only sees at 20–40% higher intensities. The impact of the heavier dopants is more variable due to their deep absorption edges, which are distinctly visible in these plots – the most extreme case is the Se-doped target, where the elevation in damage from increasing the photon energy from 12.5 keV to 13.5 keV is similar to the elevation in damage from increasing the pulse’s fluence by a factor of 3–4 (Fig. 9(a)), or from increasing the pulse’s duration by an order of magnitude (Fig. 9(b)).
@@
Inspection of Fig. 9 shows that the ionization in the doped targets consistently sees a local maximum around a LPE of 2 keV, suggesting this may serve as an approximate value for in this regime (see Sec. IV.1). Both the slow Auger electrons and fast photoelectrons ejected by light atoms are well away from this value for in typical hard X-ray SFX experiments, while the primary electrons ejected by heavy atoms are almost always closer. This is the case in the lysozyme.Gd scenario considered in Sec. III. As annotated in Fig. 3(a), the photoelectrons ejected by the heavy elements are closer to 2 keV than those of the light atoms and thus more damaging, and the same is true of the Auger electrons. The tendency of heavy atoms to eject electrons with ‘intermediate’ energies closer to thus contributes to the strength of their effect.
V Interstitial solvent
The presence of aqueous solution in targets is ubiquitous to protein SFX, encapsulating the structure and making up a substantial volume of the protein crystal [24, 78, 53]. Prior work has shown that atoms within a neutral, light-atom target will be affected by high-energy EII cascades initiated 100 nm away [29, 3]. This makes cascades instigated within the solvent a significant source of electronic damage. It becomes especially important to consider the composition of the solvent where heavy elements have a significant effect on electronic damage, as the concentrations and species of heavy elements will often be quite different between solution and protein.
In Sec. III, lysozyme.Gd was modeled as a continuous target using the density and chemical composition of only the protein. We now revisit this scenario, but model a solvated crystal using the density and chemical composition of the full unit cell of real lysozyme.Gd crystals. Approximately half of the volume is taken up by water molecules (oxygen), mixed with Gd in approximately equal concentration as in the protein [69]. We neglect the presence of any salts. Treating the electronic damage dynamics of the target as homogeneous requires the following three idealizations in place of the assumption of a continuous protein material: (I) the crystal is infinite, (II) there is instantaneous mixing of free electrons with energies above 1 keV between the interstitial solvent and protein, (III) the solvent’s emission of primary electrons below 1 keV (corresponding to light-atom Auger emissions) per unit volume is identical to the protein’s.
The results for these simulations are shown in Table 1. In each case, the damage was increased when the solvent’s contribution was accounted for. This may be attributed to the light atoms’ higher primary electron emission rate in the solvated target, due to oxygen making up the majority of the light atoms rather than more weakly absorbing carbon. Unsurprisingly, the effect of introducing Gd to both targets in equal concentrations leads to a relatively similar effect on , while sulfur’s effect is much smaller due to its reduced concentration.
The dynamics in the single particle imaging limit were also considered by modeling a target as an isolated lysozyme protein suspended within an infinite mother liquor drop. In this regime, the solution makes up essentially the entirety of the protein’s environment within the 100 nm scale spanned by the high-energy electron cascades. The high-energy cascades seeded by the protein should largely dissipate in the mother liquor, while in turn the mother liquor’s electron emissions will dominate the ionization of the protein. It is thus assumed that high-energy electrons seeded by the protein can be neglected entirely. Thus for certain targets, the ionization suffered by the protein may be modeled by considering only the solvent, which has a similar atomic density. Lysozyme.Gd suspended in pure water was not considered a viable model under this approximation, as gadolinium ions eject Auger electrons with energies on the order of 1 keV that cannot be assumed to instigate cascades well above lysozyme’s 10 nm length scale [29].
VI Discussion
@@
While it is well understood that the width of an XFEL pulse significantly affects the dose absorbed by a target (during exposure) and thus the maximum tolerable fluence [2, 29, 3, 36, 79], the results presented here suggest the absorbed dose can also be significantly affected by the presence of heavy atoms and the energy of their primary electron emissions. For all targets considered, a high-energy regime exists in which the general principle that ‘higher photon energy = less damage’ remains true. However, for certain targets, the maximum photon energies available from XFEL sources are too low to enter this regime (Fig. 9). This is a departure from conventional crystallography, where higher energies are broadly beneficial. More positively, this suggests SFX experiments may be able to control the severity of damage through the choice of X-ray frequency or solvent composition. For example, excluding gadolinium from just the interstitial solvent of the lysozyme.Gd crystal reduced by 19% for the 9 keV beam (Table 1).
@@
Elements much heavier than sulfur, such as Se and Gd, are often introduced to targets for anomalous phasing, generally in tandem with an X-ray energy just above their ionization edges [41, 42, 47, 44]. It is commonly held that difficulties in applying the technique to SFX can be mitigated by using increased heavy atom concentrations [41, 42], but the results of the present work suggest this is a trade-off in the ultrafast regime, boosting the anomalous signal at the cost of additional damage to the light atom structure. For example, Fig. 9(b) shows pulses as long as 100 fs are predicted to cause less electronic damage to regular lysozyme than a 5 fs pulse of the same fluence would to lysozyme where methionine is substituted for selenomethionine. This cost is reduced when phasing is performed with an X-ray frequency well above the absorption edge, as has been done in native phasing experiments [37, 48].
@@
There is preliminary evidence that these considerations are relevant in practice. Nass et al. attributed unusual noise in the scattering profile of lysozyme.Gd to increased radiation damage induced by Gd [27]. Our modeling supports this inference, as it found the addition of just Gd to the target (alongside C, N, O and S) increases the pump-pulse by 17–21% (see Table 1). However, Fig. 9(a) suggests that, at the fluence of this experiment (nominally ph-2, though likely lower in reality [27, 80]), the 7.1 keV photon energy used (just low enough to avoid ionizing the L3-edge of the Gd ions) would be close to the optimal choice for minimizing damage in this case. Indeed, while increasing the photon energy to 9 keV caused to fall by 32–34% for the scattering pattern of lysozyme, it rose by 9–11% for lysozyme.Gd.
The ionization of Gd ions in lysozyme.Gd under XFEL pulses has previously been investigated by Refs. [27] and [47]. The charge density of the Gd ions is inferred from the integrated electron density around the Gd site relative to a reference region of light atoms, which we refer to as the electron density ratio (EDR). Notably, this measure indicated the ionization of Gd to be unexpectedly low in both studies. Fig. 10 compares EDR values obtained from pump-probe experiments by Ref. [27] (using a light atom reference region, C20N10O10) with predictions from simulations of these experiments by AC4DC using the solvated lysozyme.Gd target. The simulated EDRs accurately reproduce the features of the experimental data, but rescaled by a factor of 2. This systematic offset may be due to AC4DC neglecting bound-to-bound electron transport – as alluded to in Sec. III.1, this was previously shown to slow the charge gain of iodine in iodobenzane [73], which contains fewer light atoms than the Gd complex. For the four EDR values where the total pulse energy has a range of 0.95–0.99 mJ (35 fs, 37 fs, 62 fs, and 112 fs), the EDR rises with increasing probe delay. We argue this is consistent with the picture of the dynamics described in Sec. III.1, which highlights that secondary ionization events – the only processes active during the dead time between pump and probe – affect the light atom dynamics far more strongly than the heavy atom dynamics. This implies that the EDR does not gauge the electron density of the heavy atoms but instead serves as a rough indicator of the ratio of secondary ionization to primary ionization. Under this perspective, the EDRs at 0 fs and 102 fs – which correspond to lower pulse energies – are raised due to the greater significance of secondary ionization at lower fluences.
Using a similar methodology to Ref. [27], Ref. [47] also attempted to measure the charge of Gd in two cases – ‘high fluence’ and ‘low fluence’, peaking at 7.8 and ph-2 respectively. The work reports only the difference in pulse-averaged Gd charge between these cases, termed the ‘charge contrast’, as measured by computing the EDR and assuming negligible light-atom ionization. A charge contrast of 8.8–12 electrons per Gd was observed; however, theoretical modeling using the XATOM toolkit predicted that the Gd charge difference ought to be 25. We repeated this modeling using AC4DC, finding an even more extreme 33.1. The apparent discrepancy between theory and experiment can be largely attributed to light atom ionization reducing the denominator of the EDR (see Fig. S4) – we obtain a charge contrast of 18.2 after making this correction. When performing this calculation, it was crucial to properly account for the cascade seeded by the Gd ions – modeling that ignored Gd ion contributions to the free electron distribution saw a value of 21.5 with this method.
VII Conclusions
The zero-dimensional non-LTE model employed in this study suggests that a significant amount of damage to biological macromolecules under XFEL illumination is seeded by heavy atoms, with even the presence of native sulfur atoms significantly affecting the damage-induced loss of coherence in the scattered wavefield. This result might appear surprising given such targets only contain heavy elements in trace quantities; however, closer inspection shows this outcome to reasonably follow from two key points: (I) Heavier species emit photo- and Auger electrons at much higher rates, considerably boosting the number of secondary ionization cascades instigated within the light atom bulk. (II) Relative to the 10 fs timescale on which the structural signal is captured, light-atom-sourced electron avalanches will either have a very low energy and thus dissipate prematurely, or have a very high energy and thus a small EII rate; in contrast, avalanches initiated by heavy atoms, with energies between these two extremes, more severely degrade the captured structural signal. The non-LTE treatment of the electron continuum was necessary to capture this latter point.
The addition of heavy atoms to the environment of proteins – such as potassium and sodium ions in the mother liquor – is routine in protein crystallography; however, the results of this work suggest that in the XFEL regime their use becomes a trade-off for additional damage. Judicious choices to reduce the number of low–intermediate energy primary electron emissions may thus improve experimental outcomes where damage is a concern, or where controlling for damage across pulse parameters is necessary. Specific to de novo refinement, anomalous phasing methodologies that allow for weaker anomalous signals would see a reduction in damage-induced noise, suggesting a strength for native phasing over artificial introduction of heavier elements such as Gd. Further, the production of primary electrons near the maximally ionizing energy can be avoided entirely with careful choice of photon frequency.
The nonlinear dynamics highlighted in this study suggest two areas for future models to incorporate. (I) The dependence of the damage on the temporal pulse profile indicates a necessity for modeling of the dynamics under more realistic SASE pulse profile statistics, for which there exist a number of contemporary approaches [81, 82, 83]. (II) The effect of heavy atoms suggests a significance to the mother liquor’s composition in conjunction with electron transfer across the crystal boundary and the interstitial solvent. It is likely, for example, that the mother liquor’s high-energy electrons replace those of the crystal to an extent dependent on the crystal’s size.
@@ VII Conclusions
The obvious limitation of the presented model is its zero-dimensional treatment. Target substructures such as metal cofactors often have order 10 nm separations, and the large-scale distribution of heavy atoms is generally non-uniform due to the differing compositions of the protein and its aqueous environment. Naively, this suggests heavy atoms produce a ‘sphere‘ of electronic damage in their local region, with the distance spanned dependent on photon energy. However, whether such heterogeneous correlations actually occur is complicated by the non-uniformity in the large-scale solvent-protein structure of real targets. A model fit for exploring this possibility will likely need to break the crystal symmetry and account for spatial variation in electron density on the global scale of the target.

### v5

@@ IV Damage by different trace elements
In this section, we present the results of simulations designed to examine how the energy-level structure of a trace, heavy element in a biological target affects electronic damage. Instead of modeling a specific protein, we consider a set of targets with densities of g cm-3 and atomic ratios of C613N193O185X10, with the ‘dopant’ X swapped out for various elements. The ionization of hydrogen was ignored for these simulations; testing showed this to be a minor approximation, increasing the average light-atom charge by 0.1 or less.
@@ IV.1 Photon and primary electron energy
Fig. 6(a) shows that the choice of dopant substantially impacts global ionization. At all energies, the intensity-averaged carbon charge varies widely across the targets. Most traces show decreasing damage with increasing photon energy, consistent with prior work [2]; however, the traces for the targets doped with Fe2+ (=26), Zn2+ (=30), and Se (=34) buck this trend, with the transition over the K-edge of the dopant increasing the average charge by roughly a factor of 2 or more in each case. This reflects the substantial contribution of the K-shell to the primary ionization of these dopants; the 1s orbitals dominate the photoabsorption cross-sections of their atoms and are replenished on a subfemtosecond timescale by Auger cycling (see Sec. III.1).
Perhaps more surprisingly, the intensity-averaged carbon charge continues to increase up to 2 above an inner-shell absorption edge, despite a decreasing photoabsorption cross-section. This is most clearly seen in Fig. 6(b). Since the separation from the edge corresponds to the energy of the photoelectrons emitted from the shell, such a result might be explained if the photoelectrons of 2 keV are unusually damaging. Thus to isolate the source of the effect, it is necessary to quantify the damage that a single primary electron emission induces. We add an artificial, low-density primary-electron peak to the continuum in the simulation of the CNO (C613N203O185) when the pulse first reaches the FWHM intensity, then track the resulting electron cascade as in Sec. III.1 (see Eq. (4))
Fig. 7 confirms that the intensity-averaged ionization caused by cascades initiated by 2 keV primary electrons is higher than those caused by 0.5 keV and by 8.5 keV primary electrons. This observation is consistent with the local maximum of ionization near the photoelectron emission energy of 2 keV in Fig. 6(b).
@@
To understand why the intermediate-energy cascade is more damaging, one can consider a free electron of energy interacting with a gas of hydrogenic atoms, with electrons bound by energy . As falls, the EII cross-section grows, causing an increased ionization rate down to [67, 76]. This creates a trade-off: higher-energy cascades progress more slowly, but have the potential to free more electrons. Indeed, this effect can be seen in comparisons of the evolution of electron cascades at various energies in neutral targets presented in prior work [3, 77]. When counting all ionization events over a fixed duration, there will be a maximally ionizing cascade energy. (Fig. 8 gives a qualitative sketch of this idea.) This implies the existence of a ‘maximally damaging’ cascade energy , loosely corresponding to the energy that maximizes the intensity-averaged ionization. Under conditions that induce significant atomic disorder during the pulse, is likely lowered by Bragg gating [78, 77, 3]. Due to the higher photoabsorption cross-section near the edge, the most damaging photons eject primary electrons with energies below ; however, simulations that control for the rate of primary ionization show the difference to be slight (about 0.5 keV, see Appendix B).
@@
Fig. 6(b) plots the average carbon charge against the lowest photoelectron energy (LPE) at emission instead of the photon energy. The traces form two distinct groups: the ‘K-group’, consisting of targets doped by an element whose ‘deepest ionizable shell’ (DIS) in the considered energy range is the K-shell (N, S, Fe, Zn, Se), and the ‘L-group’, where it is the L-shell (Ag, Xe). The split in groups again appears to be a result of Auger cycling. As the DIS of each heavy dopant is maintained near maximum occupancy by decay processes, the production rate of ionizing electrons by Ag+ and Xe8+ is scaled by a factor of 4 relative to the other dopants at an equivalent LPE, due to the higher electron capacity of their L-shells.
@@
In general, the variation in electronic damage across the targets with respect to photon frequency is predominantly controlled by the LPE of the dopants and the shell number of the DIS. However, the traces of the N, S, and Fe-doped targets deviate from the rest of the K-group in Fig. 6(b) for photon energies at or below 8 keV. In these cases, the photoelectrons ejected from the light atoms are of a low enough energy to contribute significantly to their secondary ionization (relative to the dopants). Since the Zn and Se K-edges are above 8 keV, their traces show no deviation. It is unsurprising that the L-group does not contain such outliers, as the primary electron contribution of the L-group dopants is much more significant.
@@ IV.2 Damage landscape
Contour plots of the charge accumulated by the carbon atoms, as functions of photon energy and fluence (Fig. 9(a)) or pulse width (Fig. 9(b)), show significant differences due to the dopants. The effect of sulfur on the dynamics is relatively consistent; across the parameter space of Fig. 9, the S-doped target experiences levels of ionization that the pure CNO target only sees at 20–40% higher intensities. The impact of the heavier dopants is more variable due to their deep absorption edges, which are distinctly visible in these plots – the most extreme case is the Se-doped target, where the elevation in damage from increasing the photon energy from 12.5 keV to 13.5 keV is similar to the elevation in damage from increasing the pulse fluence by a factor of 3–4 (Fig. 9(a)), or from increasing the pulse width by an order of magnitude (Fig. 9(b)).
@@
Inspection of Fig. 9 shows that the ionization in the doped targets consistently sees a local maximum around a LPE of 2 keV, suggesting this may serve as an approximate value for in this regime (see Sec. IV.1). Both the slow Auger electrons and fast photoelectrons ejected by light atoms are well away from this value for in typical hard X-ray SFX experiments, while the primary electrons ejected by heavy atoms are almost always closer. Indeed, this is the case in the lysozyme.Gd scenario considered in Sec. III. For example, the keV photoelectron peak of Cl, observable in Fig. 2(a), is closer to 2 keV than the keV light-atom photoelectrons and thus more damaging, and the same is true of the Auger electrons, which are typically keV for Cl and eV for light atoms. The tendency of heavy atoms to eject electrons with ‘intermediate’ energies closer to thus contributes to the strength of their effect.
Fig. 9 allows for the magnitude of damage in targets containing different heavy elements to be compared on the keV energy scale. This picture is only accurate in a coarse-grained sense, as resonance-related phenomena near absorption edges that modulate the fine structure of the damage landscape [79, 80, 81, 53] are not modeled here. In preparation of these diagrams, we were careful to avoid any photon energies close to resonance. We neglected ionization potential depression (IPD) [82, 83, 84, 59], which a short order of magnitude calculation shows to be limited to 50eV at biomolecular densities.
V Discussion
@@
The results presented here suggest the absorbed dose and evolution of the electronic structure are strongly affected by the presence of heavy elements. For all targets considered, a high-energy regime exists in which the general principle that ‘higher photon energy = less damage’ remains true. However, for certain targets, the maximum photon energies available from XFEL sources are too low to enter this regime (Fig. 9). This is a departure from conventional crystallography, where higher energies are broadly beneficial. More positively, this suggests SFX experiments may be able to control the severity of damage through the choice of X-ray frequency or solvent composition. For example, excluding salt from the solvent in the simulation of non-derivative lysozyme (Fig. 2) reduced the average carbon charge at fs from +1.53 to +1.28.
@@
Elements much heavier than sulfur, such as Se and Gd, are often introduced to targets for anomalous phasing, generally in tandem with an X-ray energy just above their ionization edges [42, 43, 48, 45]. It is commonly held that difficulties in applying the technique to SFX can be mitigated by using increased heavy atom concentrations [42, 43], but the results of the present work suggest this is a trade-off in the ultrafast regime, boosting the anomalous signal at the cost of additional damage to the light atom structure. For example, consider the effect of phasing with the Se K-edge on global damage. For the Se-doped target, Fig. 9(b) shows adjusting the X-ray energy from 14 keV to 12 keV is as effective at reducing the ionization of light atoms as compressing the pulse width from 50 fs to 5 fs. Alternatively, this cost might be reduced if phasing is performed with an X-ray frequency well above the absorption edge, as has been done in native phasing experiments [38, 49].
@@
There is preliminary evidence that these considerations are relevant in practice. Nass et al. attributed unusual noise in the scattering profile of lysozyme.Gd nanocrystals, absent from thaumatin nanocrystals to increased radiation damage induced by Gd [27]. Our modeling supports this inference, as it found the addition of just Gd to the target (alongside C, N, O, and S) increases the pump-pulse intensity-averaged charge of C, N, O by 25.2%, 26.9%, 26.7%, respectively. However, Fig. 9 suggests that, at the fluence of this experiment (nominally ph-2, though likely lower in reality [27, 85]), the 7.1 keV photon energy used (just low enough to avoid ionizing the L3-edge of the Gd ions) would be close to the optimal choice for minimizing damage in this case. Indeed, for the ph-2 fluence pulse simulated in Sec. III, increasing the photon energy to 9 keV caused the intensity-averaged charge of protein light atoms to rise from +0.72 to +0.74 for lysozyme.Gd in 0.1 M Gd. For reference, it fell from +0.57 to +0.34 for lysozyme in water.
@@
The obvious limitation of the presented model is its zero-dimensional treatment. Target substructures such as metal cofactors often have order 10 nm separations, and the large-scale distribution of heavy atoms is generally non-uniform due to the differing compositions of the protein and its aqueous environment. Naively, this suggests heavy atoms produce a ‘sphere’ of electronic damage in their local region, with the distance spanned dependent on photon energy. However, whether such heterogeneous correlations actually occur is complicated by the non-uniformity in the large-scale solvent-protein structure of real targets. A model fit for exploring this possibility will likely need to break the crystal symmetry and account for spatial variation in electron density on the global scale of the target. Additionally, the effect of heavy atoms suggests a significance to the mother liquor composition in conjunction with electron transfer across the crystal boundary. It is likely, for example, that the high-energy electrons originating in the mother liquor replace those of the crystal to an extent dependent on the crystal size [30].
Finally, we remark on the impact of the pulse profile on these dynamics. Previous works have generally modeled the temporal intensity profile as either square [35, 59] or Gaussian [59, 61]. The ionization dynamics proved to be sensitive to this choice (see Fig. S2). Repeating the lysozyme.Gd simulation (Fig. 2) with a square pulse of the same fluence, energy, and FWHM resulted in light-element intensity-averaged charges that were 19% lower than under the Gaussian pulse. This can be attributed to an outsized effect by the earlier EII cascades [86] induced by the Gaussian pulse during its leading tail. Such cascades have a long period to ionize the target before the bulk of the elastic X-ray scattering. This indicates a necessity for modeling of the dynamics under more realistic SASE pulse profile statistics.
VI Conclusions
The zero-dimensional non-Maxwellian model employed in this study suggests that a significant amount of damage to biological macromolecules under XFEL illumination is seeded by heavy atoms, with even the presence of native sulfur atoms significantly elevating global ionization. This result might appear surprising given such targets only contain heavy elements in trace quantities; however, closer inspection shows this outcome to reasonably follow from two key points: (I) Heavier species emit photo- and Auger electrons at much higher rates, considerably boosting the number of secondary ionization cascades instigated within the light atom bulk. (II) Relative to the 10 fs timescale on which the structural signal is captured in XFEL experiments, light-atom-sourced electron avalanches will either have a very low energy and thus dissipate prematurely, or have a very high energy and thus a small EII rate; in contrast, avalanches initiated by heavy atoms, with energies between these two extremes, more severely degrade the captured structural signal. The non-Maxwellian treatment of the electron continuum was necessary to capture this latter point.
The addition of heavy atoms to the environment of proteins – such as potassium and sodium ions in the mother liquor – is routine in protein crystallography; however, the results of this work suggest that on the short timescales of XFEL pulses their use becomes a trade-off for additional ionization. Judicious choices to reduce the number of low–intermediate energy primary electron emissions may thus improve experimental outcomes where damage is a concern, or where controlling for damage across pulse parameters is necessary. Specific to de novo refinement, anomalous phasing methodologies that allow for weaker anomalous signals would see a reduction in damage-induced noise, suggesting a strength for native phasing over artificial introduction of heavier elements such as Gd. Further, the production of primary electrons near the maximally ionizing energy can be avoided entirely with careful choice of photon frequency.